Zap Video Inc. produces two basic types of video games, Clash and Slash. Pertinent data follow (DLH = direct labor hour):

Clash Slash
Sales price (per unit) $ 400 $ 330
Costs (per unit):
Direct materials 116 80
Direct labor 85 110
Variable factory overhead (@ $15 per DLH) 110 55
Allocated fixed factory overhead (based on DLHs) 40 20
Marketing expenses (all variable) 33 40
Total costs 384 305
Operating income (per unit) $ 16 $ 25

There is insufficient labor capacity in the plant to meet the combined demand for both Clash and Slash.

Both products are produced through the same production departments.

The contribution margin per unit for Clash is:

Multiple Choice

  • $49.

  • $33.

  • $61.

  • $56.

  • $37.
Answer #1

Answer : $56

Contribution Margin Per Unit =Selling Price - total Variable cost Per Unit

For Clash

Selling Price = $ 400

Variable cost Per Unit = Direct Material +Direct labor+ Variable OH +Variable Marketing Expense

Variable Cost per Unit = 116+85+110+33 = 344

Contribution Margin Per Unit = 400-344 = $56 Per Unit